I am having major trouble with slopes. Find the slope of (-3,2) and (3,-3) every answer I have tried is not correct. I had some yesterday that I could not do that had x and y variables and I had a hard time with it as well. Any tricks or tips that may make it click with me?
Thanks a bunch!

Answer by Alan3354(69443)

Find the slope of (-3,2) and (3,-3)
--------
Slope m = diffy/diffx
m = (2 -(-3))/(-3-3) = 5/-6
m = -5/6